# Taming Throughput-Latency Tradeoff in LLM Inference with Sarathi-Serve

### v1

@@ Abstract
Each LLM serving request goes through two phases. The first is prefill which processes the entire input prompt to produce one output token and the second is decode which generates the rest of output tokens, one-at-a-time. Prefill iterations have high latency but saturate GPU compute due to parallel processing of the input prompt. In contrast, decode iterations have low latency but also low compute utilization because a decode iteration processes only a single token per request. This makes batching highly effective for decodes and consequently for overall throughput. However, batching multiple requests leads to an interleaving of prefill and decode iterations which makes it challenging to achieve both high throughput and low latency.
We introduce an efficient LLM inference scheduler Sarathi-Serve inspired by the techniques we originally proposed for optimizing throughput in Sarathi [27]. Sarathi-Serve leverages chunked-prefills from Sarathi to create stall-free schedules that can add new requests in a batch without pausing ongoing decodes. Stall-free scheduling unlocks the opportunity to improve throughput with large batch sizes while minimizing the effect of batching on latency. Our evaluation shows that Sarathi-Serve improves serving throughput within desired latency SLOs of Mistral-7B by up to 2.6$\times$ on a single A100 GPU and up to 6.9$\times$ for Falcon-180B on 8 A100 GPUs over Orca and vLLM.
@@ 1 Introduction
Figure 1: Yi-34B running on two A100 GPUs serving 128 requests from arxiv-summarisation trace.  0(a) highlights one of the many generation stalls lasting over several seconds in vLLM  [50].  0(b) shows the impact of increasing load on tail latency. Sarathi-Serve serves LLM inference with low tail latency and without generation stalls.
Optimizing throughput and latency are both important objectives in LLM inference since the former helps keep serving costs tractable while the latter is necessary to meet application requirements. In this paper, we show that current LLM serving systems are subjected to a severe tradeoff between throughput and latency. In particular, LLM inference throughput can be increased significantly with batching. However, the way existing systems batch multiple requests leads to a significant compromise on either throughput or latency e.g., 0(b) shows that increasing request load significantly increases tail latency in a state-of-the-art LLM serving system vLLM  [50].
@@
Current LLM inference schedulers can be broadly classified into two categories [^1] We classify recent schedulers Splitwise [55] and DistServe [73] under a third category “disaggregated” and discuss them in §6., namely, prefill-prioritizing and decode-prioritizing depending on how they schedule the prefill and decode phases while batching requests.
In this paper, we argue that both strategies have fundamental pitfalls making them unsuitable for serving online inference (see Figure 2).
@@
Traditional request-level batching systems such as FasterTransformer [7] employ decode-prioritizing scheduling. These systems submit a batch of requests to the execution engine that first computes the prefill phase of all requests and then schedules their decode phase. The batch completes only after all requests in it have finished their decode phase i.e., new prefills are not scheduled as long as one or more requests are doing decodes. This strategy optimizes inference for latency metric time-between-tokens or TBT – an important performance metric for LLMs. This is because new requests do not affect the execution of ongoing requests in their decode phase. However, decode-prioritizing schedulers severely compromise on throughput because even if some requests in a batch finish early, the batch execution continues until the completion of the last request.
@@
Figure 2: Current LLM serving systems involve a tradeoff between throughput and latency depending on their scheduling policy. Prioritizing prefills optimizes throughput but sacrifices TBT (time-between-tokens) tail latency whereas prioritizing decodes has the opposite effect. Sarathi-Serve serves high throughput with low TBT latency via stall-free batching.
@@
Orca [71] introduced iteration-level batching wherein requests can dynamically enter or exit a batch at the granularity of individual iterations. Iteration-level batching improves throughput by avoiding wasteful computation of request-level batching systems. Orca and several other recent systems like vLLM [21] combine iteration-level batching with prefill-prioritizing scheduling wherein they preferentially schedule the prefill phase of one or more requests first i.e., whenever GPU memory becomes available. This way, prefill-prioritizing schedulers have better throughput because computing prefills first allows subsequent decodes to operate at high batch sizes. However, prioritizing prefills leads to high latency because it pauses ongoing decodes. Since prefills can take arbitrarily long time depending on the context length of given prompts, prefill-prioritizing schedulers lead to an undesirable phenomenon we refer to as generation stalls in this paper. For example,  0(a) shows that generation stalls in vLLM can last over several seconds each.
We recently proposed Sarathi [27] – a throughput-oriented LLM inference engine based on two key ideas: chunked-prefills and decode-maximal batching. In Sarathi, chunked-prefills compute a prompt’s prefill phase over multiple iterations (each with a subset of the prompt tokens) and decode-maximal batching coalesces on-going decodes with a prefill chunk. Coalescing enables computing decode tokens in a compute-bound regime whereas chunking prompts helps in maximizing the number of decode tokens that can be coalesced with prefills. Compute iterations in Sarathi are also uniform which improves distributed inference using pipeline-parallelism by minimizing pipeline bubbles. In [27], we demonstrated how these techniques can improve inference throughput. SplitFuse [44] is also based on these techniques.
@@
Sarathi-Serve leverages Sarathi’s mechanism and improves online inference with stall-free scheduling wherein new requests join a running batch without pausing ongoing decodes. Sarathi-Serve builds upon iteration-level batching but with an important distinction: it throttles the number of prefill tokens in each iteration while admitting new requests in a running batch. This not only bounds the latency of each iteration but also makes it nearly independent of the total length of input prompts. This way, Sarathi-Serve minimizes the effect of computing new prefills on the TBT of ongoing decodes enabling both high throughput and low TBT latency.
In addition to throughput and TBT, time-to-first-byte (TTFT) is also an important performance metric for online LLM inference. TTFT denotes the latency between a request’s arrival and when its first output token appears i.e., a request’s TTFT is the sum of its scheduling delay and prefill computation time. Note that these metrics are not fully independent, i.e., increasing throughput also reduces TTFT. Sarathi-Serve enables low TBT latency and high throughput at the expense of a marginal increase in TTFT – this is due to the overhead of chunking. However, this trade-off between TBT and TTFT can be controlled by tuning the chunk size.
We evaluate Sarathi-Serve across different models and hardware — Mistral-7B on a single A100, Yi-34B on 2 A100 GPUs with 2-way tensor parallelism, LLaMA2-70B on 8 A40 GPUs, and Falcon-180B with 2-way pipeline and 4-way tensor parallelism across 8 A100 GPUs connected over commodity ethernet. For Yi-34B, Sarathi-Serve improves system serving capacity by up to 2.8$\times$ under different SLO targets. Similarly for Mistral-7B, we achieve up to 2.6$\times$ higher serving capacity. Sarathi-Serve also helps reduce the effect of pipeline bubbles on latency and throughput, resulting in up to 6.9$\times$ increase in end-to-end serving capacity for Falcon-180B.
@@ 2.1 The Transformer Architecture
Self-attention module: The self-attention module is central to the transformer architecture [64], enabling each part of a sequence to consider all previous parts for generating a contextual representation. During the computation of self-attention, first the Query ($Q$), Key ($K$) and Value ($V$) vectors corresponding to each input token is obtained via a linear transformation. Next, the attention operator computes a semantic relationship among all tokens of a sequence. This involves computing a dot-product of each $Q$ vector with $K$ vectors of all preceding tokens of the sequence, followed by a softmax operation to obtain a weight vector, which is then used to compute a weighted average of the $V$ vectors. This attention computation can be performed across multiple heads, whose outputs are combined using a linear transformation.
@@ 2.2 LLM Inference Process
Autoregressive decoding: LLM inference request processing consists of two distinct phases – a prefill phase followed by a decode phase. The prefill phase processes the entire user input prompt and produces the first output token. Subsequently, the decode phase generates output tokens one at a time wherein the token generated in the previous step is passed through the model to generate the next token until a special end-of-sequence token is generated. Note that the decode phase requires access to all the keys and values associated with all the previously processed tokens to perform the attention operation. To avoid repeated recomputation, contemporary LLM inference systems store activations in KV-cache [61, 71, 7].
@@
Figure 3: Throughput of the prefill and decode phases with different batch sizes for Mistral-7B running on a single A100 GPU. We use prompt length of 1024 for both prefill and decode experiments. Batching boosts decode throughput almost linearly but has a marginal effect on prefill throughput.
Batched LLM inference in multi-tenant environment: In production serving systems, the inference system must deal with concurrent requests from multiple users. Naively processing requests in a sequential manner leads to a severe under-utilization of GPU compute. In order to achieve higher GPU utilization, LLM inference systems leverage batching to process multiple requests concurrently. This is particularly effective for the decode phase processing which has lower computational intensity at low batch sizes. Higher batch sizes allows the cost of fetching model parameters to be amortized across multiple requests. On the contrary, prefill phase of each request contains sufficient number of token to saturate GPU compute, which renders batching ineffective. Figure 3 illustrates throughput as a function of batch size, and we can observe that while for decode iterations throughput increases almost linearly with batch size, prefill throughput almost saturates even with a single request.
Recently, several complementary techniques have been proposed to optimize throughput by enabling support for larger batch sizes. Kwon et al. propose PagedAttention [50], which allows more requests to concurrently execute, eliminating fragmentation in KV-cache. The use of Multi Query Attention (MQA) [58], Group Query Attention (GQA) [28] in leading edge LLM models like LLaMA2 [63], Falcon [29] and Yi [23] has also significantly helped in alleviating memory bottleneck in LLM inference. For instance, LLaMA2-70B model has a 8$\times$ smaller KV-cache footprint compared to LLaMA-65B.
@@
5:      while can_allocate_request($R_{new}$do
@@
2.3 Performance Metrics
Most LLM applications, e.g. ChatGPT [6], Bing AI [4], etc. follow the streaming model for request response, where tokens are streamed to the user as they are generated. In this context, there are two primary latency metrics of interest: TTFT (time-to-first-token) and TBT (time-between-tokens). For a given request, TTFT measures the latency of generating the first output token from the moment a request arrives in the system. This metric reflects the initial responsiveness of the model. TBT on the other hand measures the interval between the generation of consecutive output tokens of a request, and affects the overall perceived fluidity of the response. When system is under load, low throughput can lead to large scheduling delays and consequently higher TTFT latency.
In addition, we use a throughput metric Capacity defined as the maximum request load (queries-per-second) a system can sustain while meeting certain latency targets. Higher capacity is desirable because it reduces the cost of serving inference.
2.4 Scheduling Policies for LLM Inference
Algorithm 2 Iteration-level batching. Prefills are executed eagerly (lines 8-9), potentially introducing a generations stall for ongoing decodes (line 12).
@@ 2.4 Scheduling Policies for LLM Inference
3:   $B_{new}\leftarrow\emptyset$
4:   $R_{new}\leftarrow$ get_next_request()
5:   while can_allocate_request($R_{new}$do
@@
13:   $B\leftarrow$ filter_finished_requests($B$)
To avoid wasted compute of request-level batching, Orca [71] introduced a fine-grained iteration-level batching mechanism where requests can dynamically enter and exit a batch after each model iteration (Algorithm 2). This approach can significantly increase system throughput and is being used in many LLM inference serving systems today e.g., vLLM [21], TensorRT-LLM [20], and LightLLM [12] have adopted iteration-level batching.
Current iteration-level batching systems such as vLLM [21] and Orca [71] use prefill-prioritizing schedules that eagerly admit new requests in a running batch at the first available opportunity e.g., whenever GPU memory becomes available. Prioritizing prefills can improve throughput because it increases the batch size of subsequent decode iterations.

### v3

@@ Abstract
Each LLM serving request goes through two phases. The first is prefill which processes the entire input prompt and produces the first output token and the second is decode which generates the rest of output tokens, one-at-a-time. Prefill iterations have high latency but saturate GPU compute due to parallel processing of the input prompt. In contrast, decode iterations have low latency but also low compute utilization because a decode iteration processes only a single token per request. This makes batching highly effective for decodes and consequently for overall throughput. However, batching multiple requests leads to an interleaving of prefill and decode iterations which makes it challenging to achieve both high throughput and low latency.
We introduce an efficient LLM inference scheduler, Sarathi-Serve, to address this throughput-latency tradeoff. Sarathi-Serve introduces chunked-prefills which splits a prefill request into near equal sized chunks and creates stall-free schedules that adds new requests in a batch without pausing ongoing decodes. Stall-free scheduling unlocks the opportunity to improve throughput with large batch sizes while minimizing the effect of batching on latency. Furthermore, uniform batches in Sarathi-Serve ameliorate the imbalance between iterations, resulting in minimal pipeline bubbles.
[…]
@@ 1 Introduction
Figure 1: Yi-34B running on two A100 GPUs serving 128 requests from arxiv-summarisation trace.  1(a) highlights one of the many generation stalls lasting over several seconds in vLLM  [53].  1(b) shows the impact of increasing load on tail latency. Sarathi-Serve improves throughput while eliminating generation stalls.
Optimizing throughput and latency are both important objectives in LLM inference since the former helps keep serving costs tractable while the latter is necessary to meet application requirements. In this paper, we show that current LLM serving systems have to face a tradeoff between throughput and latency. In particular, LLM inference throughput can be increased significantly with batching. However, the way existing systems batch multiple requests leads to a compromise on either throughput or latency. For example, 1(b) shows that increasing load can significantly increase tail latency in a state-of-the-art LLM serving system vLLM  [53].
@@
Current LLM inference schedulers can be broadly classified into two categories [^1] We classify recent schedulers Splitwise [58] and DistServe [77] under a third category “disaggregated” and discuss them in §6., namely, prefill-prioritizing and decode-prioritizing depending on how they schedule the prefill and decode phases while batching requests. In this paper, we argue that both strategies have fundamental pitfalls that make them unsuitable for serving online inference (see Figure 2).
@@
Traditional request-level batching systems such as FasterTransformer [7] employ decode-prioritizing scheduling. These systems submit a batch of requests to the execution engine that first computes the prefill phase of all requests and then schedules their decode phase. The batch completes only after all requests in it have finished their decode phase i.e., new prefills are not scheduled as long as one or more requests are doing decodes. This strategy optimizes inference for latency metric time-between-tokens or TBT – an important performance metric for LLMs. This is because new requests do not affect the execution of ongoing requests in their decode phase. However, decode-prioritizing schedulers severely compromise on throughput because even if some requests in a batch finish early, the execution continues with reduced batch size until the completion of the last request.
@@
Figure 2: Current LLM serving systems involve a tradeoff between throughput and latency depending on their scheduling policy. Prioritizing prefills optimizes throughput but sacrifices TBT (time-between-tokens) tail latency whereas prioritizing decodes has the opposite effect. Sarathi-Serve serves high throughput with low TBT latency via stall-free batching. (The figure is illustrative and actual values will depend on the model and workload characteristics.)
@@
Orca [75] introduced iteration-level batching wherein requests can dynamically enter or exit a batch at the granularity of individual iterations. Iteration-level batching improves throughput by avoiding inefficiencies of request-level batching systems. Orca and several other recent systems like vLLM [23] combine iteration-level batching with prefill-prioritizing scheduling wherein they eagerly schedule the prefill phase of one or more requests first i.e., whenever GPU memory becomes available. This way, prefill-prioritizing schedulers have better throughput because computing prefills first allows subsequent decodes to operate at high batch sizes. However, prioritizing prefills leads to high latency because it interferes with ongoing decodes. Since prefills can take arbitrarily long time depending on the lengths of the given prompts, prefill-prioritizing schedulers lead to an undesirable phenomenon that we refer to as generation stalls in this paper. For example,  1(a) shows that a generation stall in vLLM can last over several seconds.
Another challenge introduced by traditional iteration-level scheduling systems like Orca [75] is pipeline stalls or bubbles [49]. These appear in pipeline-parallelism (PP) deployments that are needed to scale LLM inference across several nodes. In servers with high bandwidth connectivity such as NVIDIA DGX A100 [16], tensor-parallelism (TP) [64] can enable deployment of an LLM on up to 8 GPUs, supporting large batch sizes with low latencies. However, TP can have prohibitively high latencies when hyper-clusters are unavailable [33]. Thus, as an alternative to TP, pipeline-parallelism (PP) [55, 33] is typically used across commodity networks. Existing systems rely on micro-batches to mitigate pipeline stalls or bubbles [49]. However, the standard micro-batch based scheduling can still lead to pipeline bubbles due to the unique characteristics of LLM inference. Specifically, LLM inference consists of a mixture of varying length prefills and decodes. The resulting schedule can thus have wildly varying runtimes across different micro-batches that waste GPU cycles and degrade the overall system throughput.
@@
To address these challenges, we propose Sarathi-Serve, a scheduler to balance the throughput-latency tradeoff for scalable online LLM inference serving. Sarathi-Serve is based on two key ideas: chunked-prefills and stall-free scheduling. Chunked-prefills splits a prefill request into equal compute-sized chunks and computes a prompt’s prefill phase over multiple iterations (each with a subset of the prompt tokens). Stall-free scheduling allows new requests to join a running batch without pausing ongoing decodes. This involves constructing a batch by coalescing all the on-going decodes with one (or more) prefill chunks from new requests such that each batch reaches the pre-configured chunk size. Sarathi-Serve builds upon iteration-level batching but with an important distinction: it throttles the number of prefill tokens in each iteration while admitting new requests in a running batch. This not only bounds the latency of each iteration, but also makes it nearly independent of the total length of input prompts. This way, Sarathi-Serve minimizes the effect of computing new prefills on the TBT of ongoing decodes enabling both high throughput and low TBT latency.
In addition, hybrid batches (consisting of prefill and decode tokens) constructed by Sarathi-Serve have a near-uniform compute requirement. With pipeline-parallelism, this allows us to create balanced micro-batching based schedules that significantly reduce pipeline bubbles and improve GPU utilization, thus allowing efficient and scalable deployments.
[…]
The main contributions of our paper include:
  1. 1.
    We identify a number of pitfalls in the current LLM serving systems, particularly in the context of navigating the throughput-latency tradeoff.
  2. 2.
    We introduce two simple-yet-effective techniques, chunked-prefills and stall-free batching, to improve the performance of an LLM serving system.
  3. 3.
    We show generality through extensive evaluation over multiple models, hardware, and parallelism strategies demonstrating that Sarathi-Serve improves model serving capacity by up to an order of magnitude.
@@ 2.1 The Transformer Architecture
Self-attention module: The self-attention module is central to the transformer architecture [67], enabling each part of a sequence to consider all previous parts for generating a contextual representation. During the computation of self-attention, first the Query ($Q$), Key ($K$) and Value ($V$) vectors corresponding to each input token are obtained via a linear transformation. Next, the attention operator computes a semantic relationship among all tokens of a sequence. This involves computing a dot-product of each $Q$ vector with $K$ vectors of all preceding tokens of the sequence, followed by a softmax operation to obtain a weight vector, which is then used to compute a weighted average of the $V$ vectors. This attention computation can be performed across multiple heads, whose outputs are combined using a linear transformation.
@@ 2.2 LLM Inference Process
Autoregressive decoding: LLM inference consists of two distinct phases – a prefill phase followed by a decode phase. The prefill phase processes the user’s input prompt and produces the first output token. Subsequently, the decode phase generates output tokens one at a time wherein the token generated in the previous step is passed through the model to generate the next token until a special end-of-sequence token is generated. Note that the decode phase requires access to all the keys and values associated with all the previously processed tokens to perform the attention operation. To avoid repeated recomputation, contemporary LLM inference systems store activations in KV-cache [64, 75, 7].
@@
Batched LLM inference in multi-tenant environment: A production serving system must deal with concurrent requests from multiple users. Naively processing requests in a sequential manner leads to a severe under-utilization of GPU compute. In order to achieve higher GPU utilization, LLM serving systems leverage batching to process multiple requests concurrently. This is particularly effective for the decode phase processing which has lower computational intensity at low batch sizes. Higher batch sizes allows the cost of fetching model parameters to be amortized across multiple requests.
[…]
@@
5:  while can_allocate_request($R_{new}$) do
@@
Algorithm 2 Iteration-level batching (vLLM). Prefills are executed eagerly (lines 8-9), potentially introducing a generation stall for ongoing decodes (line 12).
@@
[…]
5:while can_allocate_request($R_{new}$) do
@@
[…]
2.3 Multi-GPU LLM Inference
With ever-increasing growth in model sizes, it becomes necessary to scale LLMs to multi-GPU or even multi-node deployments [59, 22]. Furthermore, LLM inference throughput, specifically that of the decode phase is limited by the maximum batch size we can fit on a GPU. Inference efficiency can therefore benefit from model-parallelism which allows larger batch sizes by sharding model weights across multiple GPUs. Prior work has employed both tensor-parallelism (TP) [64] and pipeline-parallelism (PP) [75, 7, 72] for this purpose.
TP shards each layer across the participating GPUs by splitting the model weights and KV-cache equally across GPU workers. This way, TP can linearly scale per-GPU batch size. However, TP involves a high communication cost due to two all-reduce operations per layer – one in attention computation and the other in FFN [64]. Moreover, since these communication operations are in the critical path, TP is preferred only within a single node where GPUs are connected via high bandwidth interconnects like NVLink.
Compared to TP, PP splits a model layer-wise, where each GPU is responsible for a subset of layers. To keep all GPUs in the ‘pipeline’ busy, micro-batching is employed. These micro-batches move along the pipeline from one stage to the next at each iteration. PP has much better compute-communication ratio compared to TP, as it only needs to send activations once for multiple layers of compute. Furthermore, PP requires communication only via point-to-point communication operations, compared to the more expensive allreduces in TP. Thus, PP is more efficient than TP when high-bandwidth interconnects are unavailable e.g., in cross-node deployments.
2.4 Performance Metrics
There are two primary latency metrics of interest for LLM serving: TTFT (time-to-first-token) and TBT (time-between-tokens). For a given request, TTFT measures the latency of generating the first output token from the moment a request arrives in the system. This metric reflects the initial responsiveness of the model. TBT on the other hand measures the interval between the generation of consecutive output tokens of a request, and affects the overall perceived fluidity of the response. When system is under load, low throughput can lead to large scheduling delays and consequently higher TTFT.
In addition, we use a throughput metric, Capacity, defined as the maximum request load (queries-per-second) a system can sustain while meeting certain latency targets. Higher capacity is desirable because it reduces the cost of serving.
2.5 Scheduling Policies for LLM Inference
To avoid wasted compute of request-level batching, Orca [75] introduced a fine-grained iteration-level batching mechanism where requests can dynamically enter and exit a batch after each model iteration.(Algorithm 2). This approach can significantly increase system throughput and is being used in many LLM inference serving systems today e.g., vLLM [23], TensorRT-LLM [21], and LightLLM [12].
Current iteration-level batching systems such as vLLM [23] and Orca [75] use prefill-prioritizing schedules that eagerly admit new requests in a running batch at the first available opportunity, e.g., whenever GPU memory becomes available. Prioritizing prefills can improve throughput because it increases the batch size of subsequent decode iterations.
@@ 3 Motivation
Figure 3: Throughput of the prefill and decode phases with different batch sizes for Mistral-7B running on a single A100 GPU. We use prompt length of 1024 for both prefill and decode experiments. Note that different y-axis, showing prefills are much more efficient than decode. Further, note that batching boosts decode throughput almost linearly but has a marginal effect on prefill throughput.
In this section, we first analyse the cost of prefill and decode operations. We then highlight the throughput-latency trade-off and pipeline bubbles that appear in serving LLMs.
3.1 Cost Analysis of Prefill and Decode
As discussed in §2.2, while the prefill phase processes all input tokens in parallel and effectively saturates GPU compute, the decode phase processes only a single token at a time and is very inefficient. Figure 3 illustrates throughput as a function of batch size, and we can observe that while for decode iterations throughput increases roughly linearly with batch size, prefill throughput almost saturates even with a single request.
Figure 4 breaks down the prefill and decode compute times into linear, attention and others, and shows their individual contributions. From the figure, we see that linear operators contribute to the majority of the runtime cost. While attention cost grows quadratically with sequence length, linear operators still contribute more than 80% to the total time even at high sequence lengths. Therefore, optimizing linear operators is important for improving LLM inference.
Low Compute Utilization during Decodes: Low compute utilization during the decode phase is a waste of GPU’s processing capacity. To understand this further, we analyze the arithmetic intensity of prefill and decode iterations. Since the majority of the time in LLM inference is spent in linear operators, we focus our analysis on them.
Matrix multiplication kernels overlap memory accesses along with computation of math operations. The total execution time of an operation can be approximated to $T=\text{max}(T_{\text{math}},T_{\text{mem}})$, where $T_{\text{math}}$ and $T_{\text{mem}}$ represent the time spent on math and memory fetch operations respectively. An operation is considered memory-bound if $T_{\text{math}}<T_{\text{mem}}$. Memory-bound operations have low Model FLOPs Utilization (MFU) [35]. On the other hand, compute-bound operations have low Model Bandwidth Utilization (MBU). When $T_{\text{math}}=T_{\text{mem}}$, both compute and memory bandwidth utilization are maximized. Arithmetic intensity quantifies the number of math operations performed per byte of data fetched from the memory. At the optimal point, the arithmetic intensity of operation matches the FLOPS-to-Bandwidth ratio of the device. Figure 5 shows arithmetic intensity as a function of the number of tokens in the batch for linear layers in LLaMA2-70B running on four A100 GPUs. Prefill batches amortize the cost of fetching weights of the linear operators from HBM memory to GPU cache over a large number of tokens, allowing it to have high arithmetic intensity. In contrast, decode batches have very low computation intensity. Figure 6 shows the total execution time of linear operators in an iteration for LLaMA2-70B as a function of the number of tokens. Note that execution time increases only marginally in the beginning i.e., as long as the batch is in a memory-bound regime, but linearly afterwards i.e., when the batch becomes compute-bound. [^2] Theoretically, we expect the operators to become compute-bound at $\sim$200 tokens on A100 GPUs, however, in practice we observe that it happens at $\sim$500-600 tokens for higher tensor parallel dimensions due to fixed overheads.
Figure 5: Arithmetic intensity trend for LLaMA2-70B linear operations with different number of token running on four A100s. Decode batches have low arithmetic intensity i.e., they are bottlenecked by memory fetch time, leading to low compute utilization. Prefill batches are compute bound with sub-optimal bandwidth utilization. Sarathi-Serve forms balanced batches by combining decodes and prefill chunks to maximize both compute and bandwidth utilization.
Takeaway-2: Decode batches operate in memory-bound regime leaving compute underutilized. This implies that more tokens can be processed along with a decode batch without significantly increasing its latency.
Figure 6: Linear layer execution time as function of number of tokens in a batch for LLaMA2-70B on A100(s) with different tensor parallel degrees. When the number of tokens is small, execution time is dictated by the cost of fetching weights from HBM memory. Hence, execution time is largely stagnant in the 128-512 tokens range, especially for higher tensor parallel degrees. Once the number of tokens in the batch cross a critical threshold, the operation become compute bound and the runtime increases linearly with number of tokens.
Figure 7: A generation stall occurs when one or more prefills are scheduled in between consecutive decode iterations of a request. A, B, C and D represent different requests. Subscript $d$ represents a decode iteration, $p$ represents a full prefill and $p0$, $p1$ represent two chunked prefills of a given prompt. vLLM induces generation stalls by scheduling as many prefills as possible before resuming ongoing decodes. Despite supporting hybrid batches, Orca cannot mitigate generation stalls because the execution time of batches containing long prompts remains high. FasterTransformer is free of generation stalls as it finishes all ongoing decodes before scheduling a new prefill but compromises on throughput due to low decode batch size. In contrast, Sarathi-Serve generates a schedule that eliminates generation stalls yet delivers high throughput.